OpenStudy (anonymous):

Copy Cats charges a $21 surcharge for aatch of colored copies plus $0.15 per page. Heritage printing charges a $25.50 surcharge for a batch of colored copies $0.12 per page. How many pages of colored copies would make Heritage Printing Cheaper than Copy Cats?

OpenStudy (ganpat):

21 + 0.15x ---- > copy cats 25.5 + 0.12x ----> heritage can u take further ??

OpenStudy (anonymous):

21/0.15=x 25.5/0.12=x

OpenStudy (ganpat):

equate 2 equations, 21 + 0.15x = 25.5 + 0.12x 0.03x = 4.5 x = 150 for 150 papers both will be charged same, so 151st paper Heritage will be cheaper

OpenStudy (anonymous):

thank you for explaining it to me

OpenStudy (ganpat):

:) W.c